### v3

# Panoramic Out-of-Distribution Segmentation

Abstract
Panoramic imaging enables capturing $360^{\circ}$ images with an ultra-wide Field-of-View (FoV) for dense omnidirectional perception, which is critical to applications, such as autonomous driving and augmented reality, etc. However, current panoramic semantic segmentation methods fail to identify outliers, and pinhole Out-of-distribution Segmentation (OoS) models perform unsatisfactorily in the panoramic domain due to pixel distortions and background clutter. To address these issues, we introduce a new task, Panoramic Out-of-distribution Segmentation (PanOoS), with the aim of achieving comprehensive and safe scene understanding. Furthermore, we propose the first solution, POS, which adapts to the characteristics of panoramic images through text-guided prompt distribution learning. Specifically, POS integrates a disentanglement strategy designed to materialize the cross-domain generalization capability of CLIP. The proposed Prompt-based Restoration Attention (PRA) optimizes semantic decoding by prompt guidance and self-adaptive correction, while Bilevel Prompt Distribution Learning (BPDL) refines the manifold of per-pixel mask embeddings via semantic prototype supervision. Besides, to compensate for the scarcity of PanOoS datasets, we establish two benchmarks: DenseOoS, which features diverse outliers in complex environments, and QuadOoS, captured by a quadruped robot with a panoramic annular lens system. Extensive experiments demonstrate superior performance of POS, with AuPRC improving by $34.25\%$ and FPR95 decreasing by $21.42\%$ on DenseOoS, outperforming state-of-the-art pinhole-OoS methods. Moreover, POS achieves leading closed-set segmentation capabilities and advances the development of panoramic understanding. Code and datasets will be available at PanOoS.
I Introduction
Omnidirectional perception technologies are vital for embodied agents, such as autonomous vehicles and mobile robots, enabling comprehensive understanding of the surroundings [zhang2024behind, ai2025survey]. In particular, panoramic imaging technology, capable of capturing high-quality $360^{\circ}$ images via optical systems, has gained significant attention across various omnidirectional vision tasks [zheng2025one_flight], such as dense visual prediction [shen2022panoformer, ling2023panoswin, yu2023panelnet], holistic scene understanding [zhang2021deeppanocontext, zheng2025open], and panoramic scene segmentation [yang2020omnisupervised, zhang2022bending, zheng2023both, cao2024occlusion]. However, existing panoramic semantic segmentation models are trained to recognize a predefined set of semantic classes (e.g., road, building, car, pedestrian[cordts2016cityscapes, ma2021densepass, huang2024high], which fail to identify outliers. Due to the broader Field-of-View (FoV) of panoramic images and the long-tail nature of semantic distribution in unconstrained surroundings, objects that do not belong to the predefined classes (e.g., animals, furniture) frequently appear in real-world driving environments, which greatly exacerbates the problem. Therefore, there is an urgent need for a technology capable of identifying outliers not present during training to avoid potential dangers.
[…]
To enhance comprehensive and safe omnidirectional scene understanding, we introduce a new task, termed Panoramic Out-of-Distribution Segmentation (PanOoS). Unlike pinhole out-of-distribution segmentation (pinhole-OoS), PanOoS offers a broader FoV beyond the pinhole view, for example, ranging from $95^{\circ}$ to $360^{\circ}$ as depicted in Fig. 1, enabling detection of outliers and even dangers not visible from a pinhole perspective, e.g., a tiger. Yet, panoramic images pose unique challenges in visual understanding tasks due to their broader FoV and complex scene content, which introduce pixel distortions [gao2022review, yang2021capturing] and background clutter, thereby complicating semantic interpretation. These characteristics lead to feature distribution shifts and loss of semantic information, impairing the model’s ability to capture low-level features and fine details of the image. Moreover, these issues interfere with foreground object recognition, introduce semantic ambiguity in overlapping or visually similar regions, and complicate distinction between in-distribution and out-of-distribution areas.
Figure 2: Comparison of panoramic semantic and out-of-distribution segmentation paradigms. (a) Panoramic semantic segmentation methods adapt the model from pinhole to panoramas through domain adaptation, yet remain limited to in-distribution dense predictions; (b) Pinhole-OoS methods combine semantic segmentation with out-of-distribution (OoD) scoring to detect outliers, but are restricted to narrow pinhole FoV; (c) POS achieves segmentation in the distribution space of panoramic images via prompt distribution learning.
Figure 3: Impact of pixel distortions and background clutter on state-of-the-art pinhole-OoS methods. Existing pinhole-OoS methods, e.g. Mask2Anomaly and RbA, fail to adapt to pixel distortions and background clutter, resulting in a higher number of false positives. In contrast, the proposed POS effectively enhances semantic perception of panoramic pixels.
Although several effective techniques exist for panoramic semantic segmentation [zhang2022bending, zheng2023look, li2023sgat4pass], they are only designed for closed-set segmentation. The inability to handle outliers means that these techniques may miss critical objects, leading to incorrect or incomplete scene understanding. This issue significantly limits their practical applicability, particularly in dynamic and unpredictable environments where outliers frequently occur. High costs associated with data annotation have also resulted in a scarcity of datasets specifically designed for training panoramic segmentation models. As illustrated in Fig. 2, most panoramic segmentation methods rely on techniques such as unsupervised domain adaptation [zheng2023both, zheng2023look, cao2024occlusion] or knowledge adaptation [zhang2024goodsam, jaus2023panoramic_insights, zhang2024goodsam++] to transfer models from the pinhole to the panoramic imagery. However, these approaches are not suitable for PanOoS, where the model must not only adapt to the characteristics of the panoramic imagery, but also learn to distinguish between in-distribution regions and outliers while ensuring closed-set segmentation performance. Meanwhile, state-of-the-art OoS methods [grcic2023advantages, nayal2023rba, rai2024mask2anomaly] for pinhole images typically involve first training a mask-based segmentation model, e.g., in Mask2Former style [cheng2022masked], and then computing anomaly scores using various outlier detection techniques during fine-tuning or inference. Nevertheless, existing research works primarily focus on narrow-FoV pinhole images, and thus lack the architectural or learning mechanisms needed to handle the geometric distortions and wide contextual dependencies inherent in panoramic imagery, making them ill-equipped for PanOoS (see Fig. 3).
To tackle the OoS challenge within the panoramic imagery, we propose POS, a mask-based framework that integrates prompt distribution learning to enable accurate pixel-wise detection of out-of-distribution regions in complex and distorted panoramic scenes. (1) Compared to purely visual models, vision-language models are typically exposed to a broader array of visual concepts during pre-training [yu2023convolutions]. To enhance the pixel-level perception of outliers within panoramic scenes, POS incorporates CLIP [radford2021learning] into a mask-transformer architecture, leveraging its strong generalization capabilities to better capture scene semantics. Moreover, during outlier exposure fine-tuning, POS explicitly disentangles per-pixel mask decoding from semantic classification, thereby mitigating semantic entanglement and preserving accurate mask representations for subsequent pixel-level decoding. (2) Traditional image-based query decoding methods [cheng2022masked, grcic2023advantages, nayal2023rba, rai2024mask2anomaly] often struggle to accurately capture semantic information under pixel distortions and background clutter. To address this, we propose Prompt-based Restoration Attention (PRA), which leverages the text encoder to extract semantic information from prompts at both class-level and distributional levels. By adaptively guiding the learning process, the PRA mechanism optimizes semantic extraction, significantly enhancing the accuracy and robustness of semantic understanding, thereby effectively mitigating semantic ambiguity in challenging panoramic scenes. (3) Finally, an innovative Bilevel Prompt Distribution Learning (BPDL) mechanism is proposed. By leveraging the language-image semantic distribution consistency of the CLIP model, BPDL refines the distribution of per-pixel mask embeddings, strengthens the distinguishability of out-of-distribution features and the adaptability of the pinhole-trained model to severe distortions encountered in panoramic images. With these crucial designs, POS achieves superior alignment with the characteristics of panoramic images, enabling efficient out-of-distribution segmentation while demonstrating impressive closed-set segmentation performance.
To facilitate the evaluation of the PanOoS task, we create two fresh datasets: (1) We have spent significant effort on artificial screening and built the first panoramic out-of-distribution segmentation dataset: Dense Panoramic Out-of-distribution Segmentation (DenseOoS). Due to the broader FoV, the backgrounds of each panoramic image are highly complex, featuring object instances that vary in shape and appearance, along with significant distortions. Meanwhile, outliers are placed in semantically plausible regions to ensure realism. (2) To further verify the performance of POS in practical application and explore the PanOoS in unconstrained, unstructured, and intense-motion real-world scenarios, we employ a quadruped mobile robot to collect panoramic images containing outliers, creating the Quadruped Panoramic Out-of-distribution Segmentation (QuadOoS) dataset. Compared to DenseOoS, QuadOoS contains a greater number of outliers per image, with smaller sizes, while maintaining similar panoramic imaging characteristics. Ultimately, based on DenseOoS and QuadOoS, we establish the evaluation benchmarks for the PanOoS task. The proposed POS achieves state-of-the-art performance on both benchmarks, significantly outperforming other pinhole-OoS methods [tian2022pixel, grcic2022densehybrid, rai2024mask2anomaly, nayal2023rba, grcic2023advantages], with AuPRC, and FPR95 of $(85.56\%,48.13\%)$, and $(0.45\%,85.22\%)$, respectively. Furthermore, compared to panoramic semantic segmentation methods [hoyer2022daformer, zhang2022bending, zhang2024behind], it demonstrates superior closed-set segmentation performance, with an mIoU of $63.30\%$ on DenseOoS. These results collectively validate the effectiveness of the proposed POS in addressing PanOoS and enhancing panoramic scene understanding.
At a glance, this work delivers the following contributions:
  • We introduce the Panoramic Out-of-Distribution Segmentation (PanOoS) task, which, for the first time, explores out-of-distribution segmentation in panoramic imagery, aiming to facilitate holistic scene understanding.
  • To achieve the PanOoS task, we propose POS, a mask-transformer-based model with Prompt-based Restoration Attention (PRA) and Bilevel Prompt Distribution Learning (BPDL), which effectively alleviates the effects of pixel distortions and background clutter, demonstrating remarkable performance.
  • Two fresh datasets, DenseOoS and QuadOoS, are created, serving as important benchmarks for PanOoS and contributing to the advancement of panoramic understanding.
II Related Work
II-A Panoramic Semantic Segmentation
By capturing broader FoV scenes, panoramic images serve as a starting point for a more holistic understanding of $360^{\circ}$ surroundings [yang2019pass, jiang2024minimalist]. The first line of research on panoramic semantic segmentation [xu2019semantic, orhan2022semantic_outdoor, li2023sgat4pass, kim2022pasts, guttikonda2024single] is based on supervised learning. However, to address the scarcity of annotations, researchers have revisited panoramic segmentation from the perspective of unsupervised domain adaptation [jang2022dada, zhu2023patch, zhang2021transfer, jiang2025gaussian, yuan2023laformer] by leveraging rich training sources from narrow-FoV pinhole data. The approaches can be divided into three types: pseudo-labeling [liu2021pano, zhang2021deeppanocontext], adversarial training [ma2021densepass, zheng2023both, jiang2025multi], and prototype adaptation [zhang2024behind, zheng2024semantics, cao2024occlusion, zheng2024360sfuda++, zhong2025omnisam]. Pseudo-labeling generates self-supervised labels for the panoramic domain. Adversarial training encourages the target model to capture domain-invariant characteristics across domains using a discriminator. Prototype adaptation approaches align high-level feature centers between the pinhole and panoramic domains. Additionally, related work demonstrates strong performance from the perspective of knowledge adaptation [zhang2024goodsam, zhang2024goodsam++, jaus2023panoramic_insights]. However, existing methods focus on closed-set segmentation and do not address the problem of identifying outliers within the scene. Recently, open panoramic segmentation [zheng2025open] has been proposed, but the focus has been on improving the efficiency of open-vocabulary segmentation. In contrast to prior work, we aim to explore out-of-distribution segmentation in panoramic images and leverage the distribution learning of text-based prompts to advance panoramic scene understanding.
II-B Out-of-distribution Segmentation
Out-of-distribution segmentation aims to identify and separate outliers in specific scenes, such as complex driving environments [chan2021segmentmeifyoucan, blum2021fishyscapes, cen2023consensus, wan2025out], while maintaining closed-set segmentation capability. Early uncertainty-based methods assume that outliers lead to low-confidence predictions and focus on estimating the prediction uncertainty through various approaches to measure pixel-wise anomaly scores [mukhoti2018evaluating, corbiere2019addressing, hendrycks2019scaling, jung2021standardized, cen2021deep]. However, models trained on closed sets may make high-confidence erroneous predictions for unseen categories. Subsequent works have shifted their focus to reconstruction. Reconstruction-based methods detect outliers by measuring the discrepancy between the original image and a reconstruction generated from semantic segmentation predictions [lis2019detecting, xia2020synthesize, haldimann2019not, vojir2021road, grcic2021dense, lis2023detecting]. However, nearly all reconstruction-based methods are highly sensitive to the quality of the reconstructed output, which can lead to performance degradation due to image artifacts [fontanel2021detecting]. These methods also suffer from context dependence and high computational overhead, resulting in unsatisfactory performance in complex environments.
Recent approaches [chan2021entropy, tian2022pixel, grcic2022densehybrid, liu2023residual, choi2023balanced] leverage the Outlier Exposure (OE) strategy [hendrycks2018deep] that utilizes auxiliary out-of-distribution data to regularize the model’s feature space. These auxiliary outliers are typically images from other datasets (such as ImageNet [deng2009imagenet], COCO [lin2014microsoft], ADE20K [zhou2017scene], etc.) or synthetic composites created by cutting out anomalous objects from external images and pasting them into inlier scenes. Notably, PEBAL [tian2022pixel] learns energy-based adaptive penalties through abstention learning, whereas DenseHybrid [grcic2022densehybrid] combines likelihood and posterior evaluation to achieve better results. Although the OE strategy generally improves performance over earlier approaches, pixel-wise architectures that adopt it still score each pixel independently without considering local semantic context, leading to noisy anomaly predictions. Currently, mask-transformer-based methods [ackermann2023maskomaly, grcic2023advantages, rai2024mask2anomaly, nayal2023rba, delic2024outlier, zhang2024csl] overcome this limitation by segmenting outliers into semantically clustered masks and encouraging the objectness of predictions, achieving leading segmentation results. EAM [grcic2023advantages] ensembles region-based outlier scores, while Mask2Anomaly [rai2024mask2anomaly] combines contrastive learning and mask refinement to significantly reduce false positives. RbA [nayal2023rba] utilizes mask queries that act as a one-vs-all classifier [grcic2023advantages, nayal2023rba, shoeb2025out], helping to eliminate uncertainty in ambiguous in-point regions, e.g., semantic boundaries, by reducing the scores in negative pixels. However, existing methods primarily focus on narrow-FoV pinhole images, and when transferred to the panoramic domain, they struggle to adapt to challenges such as pixel distortions and background clutter caused by the broad FoV. To the best of our knowledge, this is the first work to tackle this pressing problem of out-of-distribution segmentation in panoramic imagery.
III Methodology
In this work, we introduce a novel panoramic perception task, PanOoS, to promote a more comprehensive omnidirectional perception by leveraging the broader FoV of panoramic imagery. In response to the challenges posed by pixel distortions and complex backgrounds, we analyze the model architecture and fine-tuning strategies, ultimately proposing the first mask-based out-of-distribution segmentation model tailored to the PanOoS task.
III-A Problem Setting
Out-of-distribution segmentation generalizes anomaly segmentation by requiring not only the detection of anomalous pixels but also accurate semantic labeling of all in-distribution categories. Formally, given an input image $\mathbf{x}{\in}\mathbb{R}^{3\times H\times W}$, the pixel-wise segmentation architectures [cheng2022masked, nayal2023rba, rai2024mask2anomaly, grcic2023advantages, delic2024outlier] produce per-pixel class logits $\mathbf{S}(x){\in}\mathbb{R}^{K\times H\times W}$, where $K$ denotes the number of classes for closed-set segmentation, and $H$ and $W$ correspond to the height and width of the image, respectively. An anomaly score map $\mathbf{A}(x)$ is defined as
$\mathbf{A}(x)=f(\mathbf{S}(x))\in\mathbb{R}^{H\times W},$ (1)
where $f{:}~\mathbb{R}^{K}\to\mathbb{R}$ maps the per-pixel class logits to a scalar out-of-distribution confidence score.
Figure 4: Overview of POS. The proposed POS consists of four key components: an image encoder, a text encoder, a pixel decoder, and a transformer decoder with Prompt-based Restoration Attention (PRA). During outlier exposure fine-tuning, we incorporate Bilevel Prompt Distribution Learning (BPDL) to optimize the pixel manifold distribution of the per-pixel mask embeddings. See Fig. 6 for details of BPDL. During inference, the image encoder extracts multi-scale image features, which are fed into the pixel decoder to generate semantically enriched features alongside the per-pixel embeddings for segmentation. Concurrently, textual prompts are embedded through the text encoder and projected into pixel space. These prompt embeddings, along with the decoded image features, are jointly processed by the transformer decoder, where they interact iteratively with learned query features. The refined queries are subsequently used for mask prediction and category classification. Finally, an OoD scoring mechanism is applied to produce the out-of-distribution segmentation prediction.
III-B Framework
Despite the strong performance of the mask-transformer model [cheng2022masked] in pinhole-OoS, the current methods fail to generalize to PanOoS. The domain adaptation barrier prevents the model from effectively learning the visual features of panoramic images, resulting in suboptimal segmentation performance. It is important to explore methods for transferring models trained in the pinhole domain to the panoramic domain. Moreover, accurate pixel decoding is essential for panoramic segmentation. During the critical outlier exposure fine-tuning, a conflict arises between the query embedding, which favors semantic classification, and the pixel decoding, which focuses on image details. Therefore, we optimize the architecture and fine-tuning procedures, introducing POS, the first mask-transformer-based segmentation model for PanOoS, as illustrated in Fig. 4. The model comprises four main components: the image encoder, the text encoder, the pixel decoder, and the Transformer decoder.
III-B1 Image Encoder
We utilize the image encoder from CLIP [radford2021learning] for extracting image features and freeze its weights. The image encoder is a ConvNeXt-Large model pretrained in the CLIP framework [liu2022convnet], comprising four stages. It accepts an RGB image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$ as input and outputs feature maps at several resolutions to the pixel decoder. Specifically, the output feature maps are downsampled with strides of $4(\mathbf{c}_{4})$, $8(\mathbf{c}_{8})$, $16(\mathbf{c}_{16})$, and $32(\mathbf{c}_{32})$ relative to the input image.
III-B2 Text Encoder
For the textual features, the $K$ semantic class names and distribution nouns, inlier and outlier, are integrated into the pre-designed prompt templates to obtain text descriptions about classes and distributions. The text encoder is structured as a $16$-layer transformer, each layer being $768$ units wide and featuring $12$ attention heads. The descriptions are fed into the text encoder to generate the prompt embeddings $\{\mathbf{T}_{i}\mid i=0,\dots,K\}\cup\{\mathbf{P}_{in},\mathbf{P}_{out}\}{\in}\mathbb{R}^{(K+3)\times 768}$, where $\mathbf{T}_{0}$ denotes the learnable void embedding. A linear projection is then applied to project the prompt embeddings into the pixel space to match the dimensionality of per-pixel embeddings. Prompt template details are in the Appendix.
III-B3 Pixel Decoder
Following Mask2Former [cheng2022masked], the pixel decoder comprises $6$ layers of Deformable Attention Transformer (DAT) [zhu2020deformable] and a Feature Pyramid Network (FPN) [lin2017feature]. To facilitate OoS, the enriched semantic feature map $\mathbf{c}_{32}$ is processed by the DAT layers to produce $\mathbf{f}_{4}$, which is then passed to the transformer decoder. Then, the pixel decoder progressively upsamples and fuses feature maps ($\mathbf{f}_{4}$, $\mathbf{c}_{16}$, $\mathbf{c}_{8}$, and $\mathbf{c}_{4}$) to generate per-pixel mask embeddings $\mathbf{F}_{m}{\in}\mathbb{R}^{C_{m}\times H\times W}$, where $C_{m}$ is the embedding dimension.
III-B4 Transformer Decoder
The learnable query features $\mathbf{Q}{\in}\mathbb{R}^{N\times C_{q}}$ are fed into the Transformer decoder, where they are jointly processed with $\mathbf{f}_{4}$ and the prompt embeddings. $N$ and $C_{q}$ denote the number of object queries and the embedding dimension, respectively. Each decoder layer consists of a Prompt-based Restoration Attention, followed by self-attention and a Feed-Forward Network (FFN). Learnable positional embeddings are added to the query features.
III-B5 Mask and region class prediction
After refining the per-pixel embeddings and object queries, $\mathbf{Q}$ is passed through a $3$-layer MLP to obtain per-segment embeddings for mask prediction. The membership score for each pixel belonging to a region is derived by multiplying the per-segment embeddings with per-pixel embeddings $\mathbf{F}_{m}$ and applying a sigmoid function:
$\mathbf{M}(x)=\sigma(\mathrm{MLP}(\mathbf{Q})~\mathbf{F}_{m})\in\mathbb{R}^{N\times H\times W}.$ (2)
In parallel, the refined object queries are fed into a linear layer followed by a softmax function to produce posterior class probabilities $\mathbf{P}(x)$ of $K$ classes:
$\mathbf{P}(x)=\mathrm{softmax}(\mathbf{Q}\mathbf{W}_{p}^{\top})\in[0,1]^{N\times K},$ (3)
where $\mathbf{W}_{p}{\in}\mathbb{R}^{K\times C_{q}}$ is a learnable weight matrix.
III-B6 OoD Scoring
In contrast to per-pixel classification, the ground truth masks are divided into multiple binary masks, each containing all pixels belonging to a specific class. Ultimately, the class probabilities $\mathbf{P}(x)$ are broadcast to all pixels within the region, and the class logits $\mathbf{S}(x)$ are calculated as the product of the mask predictions $\mathbf{M}(x)$ with $\mathbf{P}(x)$:
$\mathbf{S}(x)=\sum_{n=1}^{N}\mathbf{P}_{n}(x)\mathbf{M}_{n}(x).$ (4)
Due to the unique architecture of mask transformer [cheng2022masked, ackermann2023maskomaly, nayal2023rba], the class logits $\mathbf{S}(x)$ aggregate weighted voting scores from all object queries; upon convergence, object queries specialize in predicting a specific class. Thus, the outlier scoring function from Rejected by All [nayal2023rba] is adopted:
$\mathbf{A}(x)=\mathrm{RbA}(\mathbf{S}(x))=-\sum_{k=1}^{K}\phi(\mathbf{S}_{k}(x)),$ (5)
with $\phi$ being the tanhtanh activation function. $\mathbf{A}(x)$ represents the outlier score map, indicating the outliers in the image $\mathbf{x}$.
III-B7 Panoramic-oriented Disentanglement
Existing research indicates that during fine-tuning, a balance must be maintained between mask segmentation and semantic prediction within the mask transformer [nayal2023rba]. We observe that PanOoS faces a similar conflict, and accurate decoding of panoramic per-pixel masks is crucial for segmentation performance under conditions of semantic clarity. Specifically, the precise decoding of foreground pixels (i.e., masks) determines the model’s capability to detect outliers. To disentangle the conflict between semantic query and per-pixel mask decoding, we introduce a disentanglement strategy specifically designed for PanOoS, and only fine-tune the pixel decoder, the $3$-layer MLP, and the linear layers of POS during the outlier exposure phase.
III-C Prompt-based Restoration Attention
Traditional query-based decoding is ineffective at capturing the intricate semantic information of panoramic imagery. This limitation primarily arises from prevalent pixel distortion and background clutter in images, which interfere with the semantic expression of queries, introduce semantic confusion, and ultimately impair the model’s comprehension of image content and segmentation performance. To address this challenge, we propose a Prompt-based Restoration Attention (PRA) module that utilizes the clear semantics embedded in text prompts to guide the semantic learning of queries. A self-adaptive mechanism that dynamically adjusts the learning process based on the semantic information of queries is proposed to enhance the robustness and generalization of semantic extraction.
Figure 5: Prompt-based Restoration Attention continuously employs masked attention and cross attention between image features and prompt embeddings while refining semantic clarity through query semantic self-adaptive correction.
Specifically, as shown in Fig. 5, we introduce an additional Cross Attention mechanism between both class and distribution prompt embeddings and the queries, built upon the existing Masked Attention. These prompt embeddings, with their clear semantic representations, provide robust semantic guidance for the query process. By incorporating the semantic information from the prompt embeddings, we mitigate potential semantic biases in the queries during the learning process, thereby significantly enhancing the semantic fidelity of queries. The interaction between the queries, image features, and prompt embeddings is formally expressed as:
$\hat{\mathbf{X}}_{i,l}$ $=\mathrm{softmax}(\mathcal{M}_{l-1}+\mathbf{Q}_{l-1}\mathbf{K}_{i,l}^{T})\mathbf{V}_{i,l},$ (6)
$\hat{\mathbf{X}}_{l}$ $=\mathrm{softmax}(\mathbf{Q}_{i}\mathbf{K}_{t}^{T})\mathbf{V}_{t},$ (7)
where $l$ denotes the layer index, and $\hat{\mathbf{X}}_{l}{\in}\mathbb{R}^{N\times C_{q}}$ represents the enriched semantic queries at the $l^{th}$ layer. $\mathbf{Q}_{l-1}$ and $\mathbf{Q}_{i}{\in}\mathbb{R}^{N\times C_{q}}$ are obtained by applying learnable linear transformations to the previous queries $\mathbf{X}_{l-1}$ and $\hat{\mathbf{X}}_{i,l}{\in}\mathbb{R}^{N\times C_{q}}$, respectively. $\mathcal{M}_{l-1}$ denotes the attention mask [cheng2022masked]. $\mathbf{K}_{i,l},\mathbf{V}_{i,l}{\in}\mathbb{R}^{H_{l}W_{l}\times C_{q}}$ are the image features under learnable transformations respectively, and $H_{l},W_{l}$ are the spatial resolution of image features. $\mathbf{K}_{t},\mathbf{V}_{t}{\in}\mathbb{R}^{(K+3)\times C_{q}}$ are the prompt embeddings extracted by the text encoder under learnable transformations, respectively.
Then, we propose a semantic self-adaptive correction mechanism, which dynamically refines the semantic representation of the query, formally expressed as:
$\mathbf{X}_{l}=\hat{\mathbf{X}}_{l}+\mathrm{FFN}(\mathrm{LN}(\hat{\mathbf{X}}_{l}))\cdot\tanh(\mathcal{W}).$ (8)
Here, LayerNorm (LN) is applied to $\hat{\mathbf{X}}_{l}$, and $\mathcal{W}$ represents the learnable adaptive weight. We initialize the weight value to zero to ensure the adaptation to diverse multi-scale image features while retaining training stability.
The proposed PRA fully leverages the clearer semantic information in prompt embeddings and incorporates an adaptive semantic correction mechanism. This not only optimizes the semantic learning process but also significantly enhances the accuracy and robustness of query semantic decoding, providing strong support for panoramic scene understanding.
III-D Bilevel Prompt Distribution Learning
Figure 6: Bilevel Prompt Distribution Learning. The distribution prompts, text prompts for each class, and pixel embeddings are denoted by stars, diamonds, and circles, respectively. The dashed circles indicate equidistant lines from the semantic center, and the dashed lines represent the decision boundary separating in- and out-of-distribution regions.
The pervasive pixel distortions and background clutter in panoramic images often result in dispersed pixel manifolds. This dispersion adversely affects the performance of deep neural networks, which tend to map the unknown samples inside the known class clusters instead of an arbitrary distribution in the whole feature space. Consequently, this misalignment distorts the voting scores of pixel manifolds, particularly in out-of-distribution regions, ultimately leading to overconfident and unreliable predictions. CLIP, trained on large-scale vision-language data pairs, demonstrates robust cross-domain generalizability. Its pretrained text encoder adeptly captures the semantic feature distribution of known classes while maintaining consistency with the pixel space distribution. This accurate semantic feature distribution provides a reliable reference for pixel manifold learning. Leveraging this foundation, we propose a Bilevel Prompt Distribution Learning (BPDL) loss that utilizes projected class and distribution prompts $\{\mathbf{T}_{i}\mid i=0,\dots,K\}\cup\{\mathbf{P}_{in},\mathbf{P}_{out}\}\in\mathbb{R}^{(K+3)\times C_{m}}$ to optimize the spatial distribution of per-pixel embeddings $\mathbf{F}_{m}$.
As shown in Fig. 6, its core objective is to enhance the discriminability of out-of-distribution pixel embeddings $\{\mathbf{O}_{i}\mid i=1,\dots,N_{o}\}\in\mathbb{R}^{N_{o}\times C_{m}}$ near the decision boundary. Specifically, the BPDL optimizes per-pixel embeddings at two levels (in- and out-of-distribution) across four key aspects.
III-D1 In-distribution learning
To improve closed-set segmentation performance, particularly by maintaining coherent intra-class representations when distinguishing out-of-distribution from in-distribution pixel embeddings $\{\mathbf{m}_{i}\mid i=1,\dots,N_{i}\}\in\mathbb{R}^{N_{i}\times C_{m}}$, we introduce an in-distribution loss $\mathcal{L}_{pixel}$. First, as shown in Fig. 7, we minimize the Euclidean distance between each pixel embedding and its corresponding class center to enhance the intra-class compactness:
$\mathcal{L}_{intra}=\frac{1}{N_{i}}\sum_{i=1}^{N_{i}}\left\|\mathbf{m}_{i}-\mathbf{T}_{k_{i}}\right\|_{2}^{2},$ (9)
where $k_{i}$ denotes the class label of the embedding $\mathbf{m}_{i}$.
Figure 7: In-distribution Learning. Enhancing intra-class compactness: the similarity of pixel embeddings within the same class is reinforced, resulting in a more compact intra-class distribution. Improving inter-class separation: the embedding distance between different classes is expanded, thereby strengthening the discriminability of inter-class features.
As demonstrated in prior works [he2018triplet, liu2022orientational], increasing the inter-class distance between category centers helps mitigate semantic ambiguity. Building upon this insight, we propose an inter-class separation loss $\mathcal{L}_{inter}$ to refine the classification boundaries in the feature space. Unlike existing approaches that rely on learned or manually defined class centers, our method leverages text prompts as predefined category prototypes, which inherently exhibit better separation due to their linguistic distinctness. We primarily propose a separation loss $\mathcal{L}_{sep}$ that increases the distance between each class center $\mathbf{T}_{i}$ and its nearest negative class center $\mathbf{T}_{n}$:
$\mathcal{L}_{sep}=\frac{1}{K+1}\sum_{i=0}^{K}\max(s-\min_{n\neq i}\left\|\mathbf{T}_{i}-\mathbf{T}_{n}\right\|_{2}^{2},\,0).$ (10)
The distance margin $s$ is defined as a hyperparameter.
In addition, to account for the spatial relationships among the remaining negative class centers in the feature space, we introduce an orientation constraint $\mathcal{L}_{ori}$ to optimize the directional relationships of class embedding distributions. As illustrated in Fig. 7 (right), when repelling the nearest negative class center $\mathbf{T}_{n}$, the angle $\theta$ between the direction $\vec{t}_{n}$ (from $\mathbf{T}_{i}$ to $\mathbf{T}_{n}$) and the direction $\vec{t}_{o}$ (from $\mathbf{T}_{i}$ to other negative class centers $\mathbf{T}_{o}$) tends to increase:
$\vec{t}_{n}=\mathbf{T}_{n}-\mathbf{T}_{i},$ (11)
$\vec{t}_{o}=\mathbf{T}_{o}-\mathbf{T}_{i},$
$\mathcal{L}_{ori}=\frac{1}{K+1}\sum_{i=0}^{K}\sum_{o\neq i,n}\omega_{o}\left(1+\cos\left(\vec{t}_{n},\vec{t}_{o}\right)\right),$ (12)
where $\omega_{o}$ represents the weights assigned to other negative class centers $\mathbf{T}_{o}$ (with $o\neq i,n$), which are determined based on their spatial relationships relative to $\mathbf{T}_{n}$:
$\omega_{o}=\frac{\exp\left(-\left\|\mathbf{T}_{n}-\mathbf{T}_{o}\right\|_{2}\right)}{\sum_{k\neq i,n}\exp\left(-\left\|\mathbf{T}_{n}-\mathbf{T}_{k}\right\|_{2}\right)}.$ (13)
The inter-class separation loss $\mathcal{L}_{inter}$ is formulated as a combination of two key terms:
$\mathcal{L}_{inter}=\mathcal{L}_{sep}+\mathcal{L}_{ori}.$ (14)
This acts as a regularization term to further constrain the feature distribution, ensuring structured separation and preventing uncontrolled dispersion of embeddings.
Ultimately, the in-distribution pixel learning loss $\mathcal{L}_{pixel}$ is defined as a sum of the intra-class compactness loss $\mathcal{L}_{intra}$ and inter-class separation loss $\mathcal{L}_{inter}$, formulated as follows:
$\mathcal{L}_{pixel}=\mathcal{L}_{intra}+\mathcal{L}_{inter}.$ (15)
Minimizing the loss $\mathcal{L}_{pixel}$ enables the model, while capturing the overall feature space distribution, to further refine its ability to differentiate between in-distribution class distributions.
Figure 8: Distribution Learning. Isolation of out-of-distribution pixels: A distinct boundary is established in the feature space to separate out-of-distribution regions from known class regions, thereby mitigating overconfident predictions on outliers. Aggregation of out-of-distribution pixels: Through semantic guidance, out-of-distribution pixel embeddings are clustered into specific semantic spaces, preventing their dispersion and interference with the feature distribution of known classes.
III-D2 Distribution learning
Building upon the aforementioned learning, a higher-level distribution learning mechanism is proposed to enhance the separability of out-of-distribution embeddings $\mathbf{O}_{i}$ and achieve a more precise delineation of the overall feature space. To establish a robust foundation for semantic alignment in subsequent distribution region partitioning and enable the model to more accurately capture the feature distribution patterns of known classes, we first minimize the distance between class text prompts $\{\mathbf{T}_{i}\mid i=0,\dots,K\}$ and in-distribution prompt $\mathbf{P}_{in}$ in the feature space:
$\mathcal{L}_{ind}=\frac{1}{K+1}\sum_{i=0}^{K}\left\|\mathbf{T}_{i}-\mathbf{P}_{in}\right\|_{2}^{2}.$ (16)
The in-distribution compactness loss $\mathcal{L}_{ind}$ promotes alignment between the feature distribution of known classes and the in-distribution semantic information.
To further constrain the distributional discrepancy between known class features and unknown pixel features, we propose a directional constraint strategy. As depicted in Fig. 8, our goal is to position known class embeddings in the opposite direction of the out-of-distribution prompt $\mathbf{P}_{out}$. Specifically, we define the direction from known class embeddings to $\mathbf{P}_{in}$ as $\vec{d}_{in}$ and the direction to $\mathbf{P}_{out}$ as $\vec{d}_{out}$. We optimize these directions to be as aligned as possible by maximizing their cosine similarity. The $\vec{d}_{in}$ and $\vec{d}_{out}$ are formulated as:
$\vec{d}_{in}$ $=\mathbf{P}_{in}-\mathbf{T}_{i},$ (17)
$\vec{d}_{out}$ $=\mathbf{P}_{out}-\mathbf{T}_{i}.$
The in-distribution directional loss $\mathcal{L}_{inDir}$ is defined as:
$\mathcal{L}_{inDir}=\frac{1}{K+1}\sum_{i=0}^{K}\left(1-\cos\left(\vec{d}_{in},\vec{d}_{out}\right)\right).$ (18)
This approach effectively mitigates confusion between in-distribution and out-of-distribution pixel embeddings in the feature space, thereby enhancing the model’s robustness to unknown features. The overall in-distribution loss $\mathcal{L}_{inlier}$ is formulated as the weighted combination of the in-distribution compactness loss $\mathcal{L}_{ind}$ and directional loss $\mathcal{L}_{inDir}$:
$\mathcal{L}_{inlier}=\alpha\,\mathcal{L}_{ind}+\mathcal{L}_{inDir},$ (19)
where $\alpha$ is the contribution weight hyperparameter of $\mathcal{L}_{ind}$.
Furthermore, to enhance the model’s ability to identify out-of-distribution regions, we employ semantic guidance to cluster out-of-distribution pixel embeddings $\{\mathbf{O}_{1},\dots,\mathbf{O}_{N_{o}}\}$ into specific regions of the semantic space. During the fine-tuning, outliers are pasted into the input images with a preset probability. This outlier exposure strategy simulates the appearance of out-of-distribution samples in real-world scenarios, enabling the model to better adapt to complex tasks in open environments. To facilitate model learning, we propose an out-of-distribution pixel loss $\mathcal{L}_{outlier}$ for both the presence and absence of outliers:
[…]
where $\{\mathbf{D}_{i,out},~\mathbf{D}_{i,in}\}=\{\|\mathbf{O}_{i}-\mathbf{P}_{out}\|_{2}^{2},~\|\mathbf{O}_{i}-\mathbf{P}_{in}\|^{2}_{2}\}$, $d$ quantifies the degree of difference between $\mathbf{P}_{in}$ and $\mathbf{P}_{out}$ within the semantic space. This approach not only preserves the stability of the known class feature distribution but also provides clear semantic attribution for out-of-distribution features. Additionally, it distinctly delineates out-of-distribution regions in the feature space, establishing a clear boundary with known class regions. As a result, it reduces the model’s overconfidence toward outliers and enhances the reliability of detection. The distribution learning loss $\mathcal{L}_{distri}$ is formulated as a sum of the in-distribution loss $\mathcal{L}_{inlier}$ and the out-of-distribution pixel loss $\mathcal{L}_{outlier}$:
$\mathcal{L}_{distri}=\mathcal{L}_{inlier}+\mathcal{L}_{outlier}.$ (21)
Finally, the BPDL loss $\mathcal{L}_{bpdl}$ is defined as follows:
$\mathcal{L}_{bpdl}=\mathcal{L}_{pixel}+\mathcal{L}_{distri}.$ (22)
Through the Bilevel Prompt Distribution Learning mechanism, the manifold of per-pixel embeddings for mask predictions becomes more compact and structured, thereby demonstrating better robustness and generalization in PanOoS.
III-E Loss Settings
Following [cheng2022masked, nayal2023rba, rai2024mask2anomaly], during training, bipartite matching is employed to match each ground truth mask with an object query, supervised by region prediction and classification losses, while the fine-tuning phase further incorporates outlier supervision and bilevel prompt distribution learning. For region prediction, a weighted combination of binary cross-entropy loss and Dice loss [milletari2016v]:
$\mathcal{L}_{mask}=\lambda_{bce}~\mathcal{L}_{bce}+\lambda_{dice}~\mathcal{L}_{dice},$ (23)
is applied to the binary mask predictions, and cross-entropy loss $\mathcal{L}_{cls}$ is used for classification to learn the semantic classes of each mask. The total training loss is defined as follows:
$\mathcal{L}=\mathcal{L}_{mask}+\lambda_{cls}~\mathcal{L}_{cls}.$ (24)
In the outlier exposure fine-tuning phase, we further utilize the $RbA$ loss $\mathcal{L}_{rba}$ [nayal2023rba] to suppress the high-confidence probabilities of outlier pixels in known classes, and the bilevel prompt distribution learning loss $\mathcal{L}_{bpdl}$ to refine the pixel manifold of per-pixel mask embeddings:
$\mathcal{L}_{oe}=\mathcal{L}+\mathcal{L}_{rba}+\lambda_{bpdl}~\mathcal{L}_{bpdl},$ (25)
where $\lambda_{bpdl}$ is the balancing coefficient.
IV Established Benchmarks
IV-A Overview of the Benchmarks
In this work, we establish two novel benchmarks: DenseOoS and QuadOoS, specifically designed for PanOoS. $360^{\circ}$ panoramas have a broader FoV and many small objects, which exaggerate the costs of creating pixel-wise annotations in unconstrained surroundings. From the perspective of the task, we aim to overcome the scarcity of PanOoS testbeds and effectively evaluate the impact of mitigating the severe distortions and background clutter on out-of-distribution segmentation performance when unmasking the narrow FoV. Furthermore, to explore PanOoS in real-world, unconstrained, unstructured, and dense-motion scenarios, we deploy a quadruped mobile robot to capture images in outdoor motion scenes with outliers, creating a real-world benchmark. Our objective with two benchmarks is to provide a comprehensive evaluation of methods capable of performing PanOoS. For comparison with previous pinhole-OoS methods, we adopt Area under PRC Curve (AuPRC/AP) and False Positive Rate at a true positive rate of $95\%$ (FPR95) for panoramic out-of-distribution segmentation pixel-wise evaluation, and mean Intersection-over-Union (mIoU) for closed-set segmentation evaluation. We unfold a comprehensive explanation in Appendix B-B.
Datasets Year Domain Annotation OoD realism No shift
Road Anomaly [lis2019detecting] 2019 Pinhole outlier
Fishscapes L&F [blum2021fishyscapes] 2019 outlier (✓)
Fishscapes Static [blum2021fishyscapes] 2019 outlier
StreetHazards [hendrycks2019scaling] 2019 all (✓)
SMIYC RA-21 [chan2021segmentmeifyoucan] 2021 outlier
SMIYC RO-21 [chan2021segmentmeifyoucan] 2021 outlier
DenseOoS 2025 Panoramic all (✓) (✓)
QuadOoS 2025 outlier
TABLE I: Comparison of different OoS datasets. Referring to POC [de2024placing], we qualitatively compare datasets on two main axes, and score them as good (), medium ((✓)) or bad ().
IV-B DenseOoS
We introduce a novel test dataset: Dense Panoramic Out-of-Distribution Segmentation (DenseOoS) tailored for PanOoS. DenseOoS comprises a labeled test set of $1,000$ panoramic images with resolution $2048{\times}400$ and provides pixel-wise annotations for the PanOoS task, which greatly extend the semantic labels from DensePASS [ma2021densepass]. Relevant studies have shown that using generated outlier data for model training or evaluation is an effective approach [loiseau2025reliability, de2024placing]. Therefore, based on the DensePASS dataset, we use a state-of-the-art image generation technique, POC [de2024placing], to place multiple classes of outliers into images while generating annotation labels for these outliers to construct DenseOoS.
The normal labels cover $19$ classes that align with those of Cityscapes [cordts2016cityscapes], while the abnormal labels contain $30$ classes of outliers as demonstrated in Fig. 9. To ensure the high quality and plausibility of the generated images, particularly that the outliers are generated in realistic areas within the real street backgrounds, our screening process follows the “preliminary screening $\rightarrow$} independent screening $\rightarrow$} voting” workflow, eliminating unreasonable object locations, and normal situations. Specifically, a total of $14,031$ images are generated, of which $1,244$ are retained after preliminary manual screening. Following an independent screening and voting process by three examiners, we ultimately select $1,000$ images for the construction of the dataset. The outliers are diverse and distributed within contextually plausible regions of the images. Importantly, due to the $360^{\circ}$ panoramic layout, over $68\%$ of these outliers lie outside the typical $95^{\circ}$ narrow-FoV of pinhole cameras, which makes DenseOoS particularly suitable for the evaluation of the PanOoS task. More details and qualitative examples are provided in the Appendix.
Figure 9: Distribution details of DenseOoS. DenseOoS contains $30$ distinct outliers, divided into 55 main categories: Animals, Garbage, Furniture, Vehicles & Tools, and Others.
IV-C QuadOoS
To further explore the challenges of the PanOoS task in real-world applications, we collect panoramic images featuring outliers in unconstrained, unstructured, and dense-motion real scenes, and manually annotate these outliers to establish the first Quadruped Panoramic Out-of-distribution Segmentation dataset, QuadOoS. This dataset provides a crucial foundation for panoramic out-of-distribution segmentation research and facilitates the evaluation and optimization of PanOoS models in complex environments. We develop a mobile platform using a quadruped robot, and the robot is selected for its biomimetic gait, a complex form of locomotion found in nature, which introduces additional challenges for PanOoS. It can navigate vertical obstacles up to $15cm$ and inclines up to $30^{\circ}$, making it highly maneuverable in everyday environments. With $12$ joint motors, the robot replicates realistic walking motions at speeds up to $2.5m/s$. For sensing, we use a customized panoramic annular camera to capture a broader FoV ($360^{\circ}{\times}70^{\circ}$). The camera has a pixel size of $3.45{\mu}m{\times}3.45{\mu}m$, a resolution of $5$ million effective pixels. Mounted on the quadruped robot (see Fig. 10(a)), the camera ensures an optimal field of view.
Using this platform, we conduct data collection in campus environments with outliers, creating the Quadruped Panoramic Out-of-distribution Segmentation (QuadOoS) dataset. In particular, we collect data across $8$ blocks in and around the campus, with each block averaging approximately $50m$ in length. The collected raw images consist of a total of $4,800$ frames ($600$ frames per block), each with an original size of $2048{\times}476$ pixels. To avoid misleading the model due to overexposed regions at the center of the lens and to maintain consistency with DenseOoS, we crop the upper portion of the original images (i.e., the sky), which does not affect fair evaluation. Subsequently, to ensure the diversity of outliers and the satisfactory imaging quality, we retain a final selection of $106$ original panoramic images after manual screening by five examiners, thereby constructing the QuadOoS dataset. Detailed information is provided in Fig. 10(b); see Appendix B-A for details of outlier types and examples from QuadOoS.
Due to the gait of the quadruped robot, the collected panoramic images exhibit irregular shaking, particularly along the Y-axis (see Fig. 10(c)). Compared to panoramic images captured under static conditions, the images in QuadOoS present significant challenges to existing methods, such as uneven exposure, color inconsistencies due to the broader FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect. Each image contains an average of $7$ outliers, accurately reflecting real-world situations and posing substantial challenges for PanOoS methods. As illustrated in Table I, with a broader FoV ($360^{\circ}{\times}70^{\circ}$), QuadOoS significantly differs from traditional pinhole-OoS datasets [lis2019detecting, blum2021fishyscapes, hendrycks2019scaling, chan2021segmentmeifyoucan], and complements DenseOoS by enabling the exploration of motion-induced and panoramic-specific challenges in PanOoS.
((a))
Frames Resolution Locations Objects Annotations
Test 106 $2048{\times}400$ 8 $\geq$18 735
((b))
((c))
Figure 10: (a) Quadruped robot platform with a panoramic annular camera; (b) is the data statistics; (c) Y-axis pixel positions of quadruped robot motion trajectories. While in motion, the quadruped robot experiences significant vertical jitter.
V Experiment Results and Analysis
V-A Datasets
We train the model on Cityscapes [cordts2016cityscapes] and an extra inlier dataset: Mapillary Vistas [neuhold2017mapillary] for outlier exposure fine-tuning. For PanOoS evaluation, we use DenseOoS and QuadOoS benchmarks. Additionally, we use Road Anomaly [lis2019detecting] to evaluate pinhole-OoS.
  • Cityscapes consists of $2,975$ training and $500$ validation images and contains $19$ categories that are considered inliers in out-of-distribution segmentation benchmarks, such as road, sky, building, and car.
  • Mapillary Vistas is a large-scale street-level image dataset containing $25,000$ high-resolution images with a variety of weather, seasons, times of day, cameras, and viewpoints, providing rich visual information.
  • Road Anomaly is a collection of $60$ web images with diverse anomalous objects on or near the road.
V-B Implementation Details
V-B1 Closed-Set Training
We follow [cheng2022masked] and adopt the same training recipe and losses without any special design. The difference is that we adopt the ConvNeXt-Large CLIP [liu2022convnet] from OpenCLIP [ilharco_gabriel_2021_5143773], pre-trained on the LAION-2B dataset [schuhmann2022laionb], and freeze its weights. In reference to previous research works [rai2024mask2anomaly, nayal2023rba], only one decoder layer with $100$ queries is used in the transformer decoder. Both dimensions $C_{m}$ and $C_{q}$ are set to $256$. The model is trained for $90k$ iterations using a batch size of $16$ on $4$ NVIDIA GeForce RTX 3090 GPUs. The training is optimized with the AdamW optimizer and a weight decay $0.05$. We employ the initial learning rate of $1{\times}10^{-4}$ and a polynomial schedule.
V-B2 Outlier Exposure Fine-tuning
After training in the closed-set setting, we fine-tune the pixel decoder, the $3$-layer MLP, and the linear layers for $30k$ iterations on the inlier set of Cityscapes [cordts2016cityscapes] and Mapillary Vistas [neuhold2017mapillary], using the setting of closed-set training; AdamW optimizer with $0.05$ weight decay and $1{\times}10^{-4}$ initial learning rate with polynomial schedule. For each inlier image used in fine-tuning, an object from the MS-COCO samples [lin2014microsoft] is uniformly chosen as an outlier and pasted on the image using AnomalyMix [tian2022pixel] with a per-image probability $p_{out}$ (set to $0.3$).
V-C Results of Panoramic Out-of-distribution Segmentation
We deliver PanOoS results on the DenseOoS and QuadOoS benchmarks. Two representative types of existing state-of-the-art pinhole-OoS methods, pixel-wise architectures, and mask-transformer-based methods, are benchmarked. As shown in Table II, mask-transformer-based OoS methods demonstrate strong potential. Compared to the best-performing method RbA [nayal2023rba], POS outstrips it by $34.25\%$ in AuPRC, $21.42\%$ in FPR95, and $5.76\%$ in mIoU on DenseOoS. Furthermore, state-of-the-art panoramic semantic segmentation models [hoyer2022daformer, zhang2022bending, zhang2024behind, zheng2024360sfuda++] are compared (see Table III). Existing panoramic semantic segmentation methods cannot distinguish between in- and out-of-distribution regions and are hindered by outliers, resulting in degraded closed-set segmentation performance. However, POS achieves excellent out-of-distribution segmentation while maintaining outstanding closed-set segmentation performance, demonstrating the strong potential of prompt distribution learning in omnidirectional scene perception. We further explore PanOoS in unconstrained, unstructured, and dense-motion real-world scenarios, where all baseline methods struggle to make accurate predictions under motion blur, resulting in significant false positive issues. Specifically, pixel-wise methods are not very sensitive to blur but struggle to effectively recognize outliers within the scene, while mask-transformer-based methods tend to produce false positives. POS achieves a good balance, with mean AuPRC and FPR95 improvements of approximately $24\%$ and $13\%$, respectively, on the two benchmarks. Benefiting from the generalization ability of vision-language models, Prompt-based Restoration Attention (PRA) and Bilevel Prompt Distribution Learning (BPDL), POS performs excellently on the two benchmarks for PanOoS, perfectly adapting to the background clutter and pixel distortions caused by the broader FoV of panoramic images.
Method DenseOoS QuadOoS Mean
AuPRC $\uparrow$ FPR95 $\downarrow$ mIoU $\uparrow$ AuPRC $\uparrow$ FPR95 $\downarrow$ AuPRC $\uparrow$ FPR95 $\downarrow$
PEBAL [tian2022pixel] 16.39 7.59 46.29 11.34 39.25 13.87 23.42
DenseHybrid [grcic2022densehybrid] 24.44 7.74 45.19 19.09 33.86 21.77 20.80
EAM [grcic2023advantages] 17.97 6.46 61.58 33.05 79.66 25.51 43.06
Mask2Anomaly [rai2024mask2anomaly] 36.28 8.79 48.10 25.93 92.55 31.10 50.67
RbA [nayal2023rba] 51.31 21.87 57.54 34.12 90.34 42.71 56.11
Ours 85.56 0.45 63.30 48.13 85.22 66.85 42.84
TABLE II: Panoramic out-of-distribution segmentation quantitative evaluation. POS shows significant improvement over baseline per-pixel and mask-transformer-based pinhole-OoS methods on PanOoS. Higher values for AuPRC and mIoU are better, whereas for FPR95 lower values are better. The best and second best results are bold and underlined, respectively, and the line in-between the table divides the per-pixel architectures and mask-transformer-based methods.
Method mIoU
DAFormer [hoyer2022daformer] 44.53
Trans4PASS (Tiny) [zhang2022bending] 44.93
Trans4PASS (Small) [zhang2022bending] 49.35
Trans4PASS+ (Tiny) [zhang2024behind] 49.29
Trans4PASS+ (Small) [zhang2024behind] 50.11
360SFUDA++ [zheng2024360sfuda++] 50.46
Ours 63.30
TABLE III: Quantitative closed-set segmentation results on DenseOoS. POS achieves outstanding panoramic semantic segmentation performance.
Method AuPRC FPR95
PEBAL [tian2022pixel] 62.37 28.29
DenseHybrid [grcic2022densehybrid] - -
EAM [grcic2023advantages] 69.40 7.70
Mask2Anomaly [rai2024mask2anomaly] 79.70 13.45
RbA [nayal2023rba] 85.42 6.92
Ours 89.39 6.25
TABLE IV: Pinhole out-of-distribution segmentation results on Road Anomaly. The best and second best results are bold and underlined, respectively; “-” indicates the unavailability of benchmark results.
V-D Results of Pinhole Out-of-distribution Segmentation
To further investigate the segmentation capacity of the proposed POS in the pinhole domain, we evaluate its OoS performance on one standard pinhole-OoS datasets [lis2019detecting]. During the outlier exposure fine-tuning, we do not apply panoramic-oriented disentanglement and fine-tune only the MLP and the linear layers. The results in Table IV indicate that POS excels on Road Anomaly, and achieves precise segmentation of outliers with minimal false positives.
V-E Ablation Study
In this section, we conduct comprehensive ablation experiments to validate the effectiveness of our solution. To ensure a rigorous comparison, all the results reported in this section are based on the DenseOoS benchmark.
V-E1 Each Component in POS
Table V presents the component-wise ablation results of the technical novelties included in POS. Using the image encoder alone as the backbone, the model demonstrates strong potential for omnidirectional visual perception. However, it still struggles with semantic ambiguity and background interference in complex panoramic scenes. By incorporating Prompt-based Restoration Attention (PRA), the model significantly improves pixel-level semantic associations through prompt distribution learning, enhancing panoramic perception. This results in an increase in AuPRC to $83.94\%$ ($7.06\%$) and an improvement in mIoU to $62.50\%$ ($0.77\%$), highlighting the effectiveness in mitigating dense pixel semantic blurring. Further incorporation of Bilevel Prompt Distribution Learning (BPDL) leads to a synergistic optimization: PRA enhances inter-class discrimination through semantic decoupling, while BPDL preserves topological consistency via manifold learning. This combined mechanism enables the model to achieve groundbreaking performance in the PanOoS task: $85.56\%$ AuPRC, $0.45\%$ FPR95, and $63.30\%$ mIoU. Not only does it overcome the coupling challenges of semantic ambiguity and geometric distortion that traditional pinhole-OoS methods cannot deal with, but it also establishes a new benchmark for panoramic scene understanding.
PRA BPDL AuPRC FPR95 mIoU
76.88 0.84 61.73
83.94 0.46 62.50
85.56 0.45 63.30
TABLE V: Component-wise ablation study of POS. PRA: Prompt-based Restoration Attention; BPDL: Bilevel Prompt Distribution Learning.
MA CA SAC AuPRC FPR95 mIoU
81.24 0.67 58.03
81.98 0.53 62.48
85.56 0.45 63.30
TABLE VI: Structural ablation of Prompt-based Restoration Attention. MA: Masked Attention; CA: Cross Attention; SAC: Self-adaptive Correction.
((a))
((b))
((c))
Figure 11: Attention maps of queries in PRA after different query stages, resembling a confusion matrix in which each query has the highest semantic similarity with itself. When querying only with image features, the semantic information of queries is relatively vague. However, after further interaction with the prompt embeddings, the semantics become clearer, and the self-adaptive correction further mitigates the impact of pixel distortion on semantic decoding.
V-E2 Prompt-base Restoration Attention
To better understand the effect of Prompt-based Restoration Attention (PRA), we perform an ablation analysis of its structure in Table VI. Experimental results show that using Masked Attention (MA) alone, which interacts only with image features, cannot adapt to the characteristics of panoramic images, resulting in poor performance across almost all metrics. This demonstrates that prompt-based Cross Attention (CA) effectively captures high-level semantic interactions, alleviating the impact of pixel distortions on semantic decoding. Furthermore, to prevent blind confidence during the learning process, the proposed Self-adaptive Correction (SAC) dynamically adjusts the semantic confidence for each query, yielding more accurate semantic segmentation and avoiding excessively high voting scores for out-of-distribution areas. This is visually substantiated in Fig. 11, where the inter-query attention maps after three stages explicitly manifest the effectiveness of PRA for hierarchical semantic refinement.
$\mathcal{L}_{pixel}$ $\mathcal{L}_{ind}$ $\mathcal{L}_{inDir}$ $\mathcal{L}_{outlier}$ AuPRC FPR95 mIoU
81.72 0.54 62.69
82.61 0.50 62.67
83.61 0.49 63.04
83.37 0.49 63.06
83.08 0.50 62.89
85.56 0.45 63.30
TABLE VII: Ablation study of different loss contributions in bilevel prompt distribution learning loss $\mathcal{L}_{bpdl}$.
((a))
((b))
((c))
((d))
Figure 12: Sensitivity analysis of $s$, $d$, $\alpha$ and $\lambda_{bpdl}$.
V-E3 Bilevel Prompt Distribution Learning
The proposed BPDL belongs to the domain of metric learning. Unlike traditional methods that rely on manually defined or learned centroid embeddings as class prototypes, BPDL uses semantically rich text embeddings as prototypes. This approach allows for a coarse-grained predefinition of distribution regions for each class. Moreover, BPDL utilizes a bilevel learning framework for distribution optimization: the lower-level optimization refines the inter-class distributions for known classes, while the upper-level optimization adjusts the global distribution space. As shown in Table VII, each component of the $\mathcal{L}_{bpdl}$ contributes to feature distribution optimization, with the $\mathcal{L}_{ind}$ and $\mathcal{L}_{pixel}$ loss terms leading to particularly notable performance gains (AuPRC: $2.95\%/{+}2.48\%$, mIoU: $0.63\%/{+}0.41\%$). Ablation experiments confirm that integrating all loss components results in optimal performance across all evaluation metrics.
As outlined in Eqs. 10, 19, 20, and 25, the BPDL framework involves four key hyperparameters: inter-class margin $s$, distribution margin $d$, loss weight of $\mathcal{L}_{Dis}$ $\alpha$, and balancing coefficient $\lambda_{bpdl}$. The inter-class margin $s$ controls the trade-off between intra-class embedding compactness and inter-class separability, while the distribution margin $d$ governs the geometric structure of the embedding space. The parameter $\alpha$ adjusts the relative contribution of the $\mathcal{L}_{Dis}$ loss, and $\lambda_{bpdl}$ balances the weights between the $\mathcal{L}_{bpdl}$ loss and other loss terms. Through a systematic sensitivity analysis (Fig. 12), we observe a notable convex trend for $s$. Both mIoU and AuPRC exhibit convex behavior as functions of $s$, while FPR95 shows a concave trend but remains stable within a fluctuation range of $0.03\%$. Similarly, the distribution margin $d$ follows a similar optimization trend, confirming the synergistic effect of the bilevel distribution constraints. The parameter $\alpha$ regulates the distance between the upper-level class prototype embeddings and the distribution embeddings, enhancing the isolation of out-of-distribution embeddings. However, an excessive emphasis on this constraint can reduce closed-set semantic segmentation performance. Additionally, when $\lambda_{bpdl}$ exceeds a certain threshold ($0.01$), the model tends to focus too much on the out-of-distribution detection task, leading to significant degradation in closed-set segmentation performance (with mIoU decreasing by $0.6\%$), while out-of-distribution detection metrics remain relatively stable.
VI Conclusion
In this work, we introduce a novel task, termed Panoramic Out-of-distribution Segmentation (PanOoS), which aims to enhance panoramic holistic scene understanding. To address PanOoS, we propose the first panoramic out-of-distribution segmentation solution, POS, which effectively tackles the challenges posed by panoramas, including pixel distortions and background clutter. The proposed PRA optimizes the interaction between visual features and text embeddings, significantly improving semantic understanding in complex panoramic scenes and mitigating semantic confusion. The BPDL further strengthens the model’s ability to recognize out-of-distribution regions by optimizing the embedding space in a hierarchical manner. The synergistic combination of PRA and BPDL allows POS to effectively resolve issues related to semantic ambiguity and geometric distortion in panoramic images. Moreover, we establish DenseOoS and QuadOoS benchmarks to facilitate the optimization and evaluation of PanOoS models, paving the way for future research in PanOoS. Extensive experiments on these two benchmarks demonstrate that POS achieves state-of-the-art performance in PanOoS, significantly improving segmentation accuracy, robustness, and the detection of out-of-distribution regions.
Appendix A Methodology
A-A Prompt Templates
We adopt the prompt templates from [radford2021learning, yu2023convolutions], but further optimize them to more accurately describe items from different distributions. Specifically, the class and distribution prompt templates are as follows:
A-A1 Class
“A photo of a {}.”
“This is a photo of a
“There is a
{} in the scene.”
“There is the
{} in the scene.”
“A photo of a
{} in the scene.”
“A photo of a small
“A photo of a medium
“A photo of a large
“This is a photo of a small
“This is a photo of a medium
“This is a photo of a large
“There is a small
{} in the scene.”
“There is a medium
{} in the scene.”
“There is a large
{} in the scene.”
A-A2 Distribution
“A photo of an {} item.”
“This is a photo of an
{} item.”
“There is an
{} item in the urban or rural scene.”
“There is the
{} item in the urban or rural scene.”
“A photo of an
{} item in the urban or rural scene.”
“A photo of a small
{} item.”
“A photo of a medium
{} item.”
“A photo of a large
{} item.”
“This is a photo of a small
{} item.”
“This is a photo of a medium
{} item.”
“This is a photo of a large
{} item.”
“There is a small
{} item in the urban or rural scene.”
“There is a medium
{} item in the urban or rural scene.”
“There is a large
{} item in the urban or rural scene.”
Appendix B Benchmarks and Experiments
B-A More Details of Datasets
B-A1 DenseOoS
Class ID Number
stroller 19 19
trolley 20 20
garbage bag 21 55
wheelie bin 22 41
suitcase 23 16
skateboard 24 13
chair dumped on the street 25 23
sofa dumped on the street 26 41
furniture dumped on the street 27 32
matress dumped on the street 28 63
garbage dumped on the street 29 90
clothes dumped on the street 30 101
cement mixer on the street 31 48
cat 32 10
dog 33 10
bird flying 34 10
horse 35 34
skunk 36 38
sheep 37 48
crocodile 38 19
alligator 39 12
bear 40 65
llama 41 18
tiger 42 38
monkey 43 7
cattle 44 43
duck 45 28
cock 46 12
box 47 25
rock 48 21
total - 1000
TABLE B.1: Details of the DenseOoS dataset.
Figure B.1: Examples from the DenseOoS dataset. Image backgrounds are highly cluttered, and distortions are severe. Outliers appear in semantically plausible regions. Zoom in for a better view.
The specific types of outliers in DenseOoS are listed in Table B.1. The normal labels cover $19$ classes that align with those of Cityscapes [cordts2016cityscapes], while the abnormal labels contain $30$ classes of outliers, including animals, garbage, and furniture. DenseOoS additionally includes $5$ other common anomalous objects (cattle, duck, cock, box, and rock) compared to POC [de2024placing]. To ensure the high quality and rationality of the generated images, particularly that the outliers are generated in realistic areas within the real street scenes, our screening process follows the “preliminary screening $\rightarrow$} independent screening $\rightarrow$} voting” workflow, eliminating unreasonable object locations and normal situations, e.g., birds flying in the city sky or garbage/garbage bags/wheelie bin placed on the sidewalk.
The overall process of dataset generation and construction is as follows. First, all text prompts for image inpainting (a good photo of a $\{class\}$) are determined based on the DensePASS validation dataset. Next, we select appropriate regions to be inserted (e.g., the road and sidewalk) and the selected rectangular area is scaled to match the resolution of the real panoramic images. Then, we set different random seeds and use the POC algorithm to batch-generate images and outlier annotations. Finally, the suitable images for constructing the DenseOoS dataset are determined through preliminary manual screening, independent screening, and majority voting. Specifically, a total of $14,031$ images are generated, of which $1,244$ are retained after preliminary manual screening. Following an independent screening and voting process by three examiners, we ultimately select $1,000$ images for the construction of DenseOoS. As shown in Fig. B.1, the outliers are diverse and distributed within semantically plausible regions of the images. A significant portion of these outliers extends beyond the pinhole narrow FoV, making them particularly suitable for the evaluation of the PanOoS task.
B-A2 QuadOoS
In addition, we list some types of outliers in QuadOoS. As presented in Fig. B.2, the outliers mainly consist of two parts: human-defined objects and outliers existing in the neighborhood. Compared to panoramic images captured under static conditions, the images in QuadOoS (see Fig. B.3) pose significant challenges to existing methods, such as uneven exposure, color inconsistencies due to the broader FoV, and increased motion blur, as rapid relative displacement between moving objects and the background intensifies the blurring effect. Each image contains an average of $7$ outliers, accurately reflecting real-world situations and posing substantial challenges for PanOoS methods.
Figure B.2: Examples of outliers included in the QuadOoS dataset. Including but not limited to the above types.
Figure B.3: Examples from the established QuadOoS dataset. The images come with challenges, including uneven exposure, color inconsistencies due to the broader FoV, and increased motion blur. Numerous small outliers are scattered across the scene.
B-B Evaluation Metrics
In the context of the Panoramic Out-of-distribution Segmentation (PanOoS) benchmark, we employ three metrics, namely Area under PRC Curve (AuPRC), False Positive Rate at a true positive rate of $95\%$ (FPR95), and mean Intersection-over-Union (mIoU), to evaluate the performance of models. For pixel-wise evaluation, $\mathcal{G}{\in}\{\mathcal{G}_{in},\mathcal{G}_{out}\}$ is the annotated ground truth labels for an image containing outliers, where $\mathcal{G}_{in}$ and $\mathcal{G}_{out}$ represent the inlier and outlier labels, respectively. Let $\mathcal{G}_{p}$ denote the model prediction obtained by a threshold $\tau$. The True Positive Rate (TPR), precision, and recall can be written as:
$TPR(\tau)\!=\!recall(\tau)\!$ $=\!\frac{\mid\!\mathcal{G}_{p}\!\cap\!\mathcal{G}_{out}\!\mid}{\mid\!\mathcal{G}_{out}\!\mid},$ (26)
$precision(\tau)\!$ $=\!\frac{\mid\!\mathcal{G}_{p}\!\cap\!\mathcal{G}_{out}\!\mid}{\mid\!\mathcal{G}_{p}\!\mid}.$ (27)
The AuPRC can be approximated as:
$AuPRC\!$ $=\!\int_{\tau}\!precision(\tau)\,recall(\tau).$ (28)
The AuPRC works well for unbalanced datasets, which makes it particularly suitable for out-of-distribution segmentation since all the datasets are significantly skewed. Next, we consider the FPR95, an important criterion for safety-critical applications, which is calculated as:
$FPR_{95}\!=\!FPR(\tau_{95})\!=\!\frac{\mid\!\mathcal{G}_{p}(\tau_{95})\!\cap\!\mathcal{G}_{in}\!\mid}{\mid\!\mathcal{G}_{in}\!\mid},$ (29)
where $\tau_{95}$ is the threshold at which the TPR equals 95%.
To effectively evaluate the in-distribution segmentation performance of the model, the standard mIoU is used for performance evaluation. The Intersection-over-Union (IoU) measures the overlap between predicted semantic segment $p$ and ground truth semantic segment $g$, while mIoU is the mean of IoU for each class in the dataset. It is calculated as:
$mIoU\!=\frac{1}{K}\!\sum_{k=1}^{K}\!\frac{p_{k}\!\cap\!g_{k}}{p_{k}\!\cup\!g_{k}}.$ (30)
B-C Comparative Experiments
For the experiments on PanOoS, we compare POS to the state-of-the-art pinhole-OoS methods: per-pixel architectures [tian2022pixel, grcic2022densehybrid] and mask-transformer-based [grcic2023advantages, rai2024mask2anomaly, nayal2023rba] methods, using the models that were shared in their respective repositories.
Appendix C More Quantitative Results
C-A More in-distribution segmentation results
Method road sidewalk building wall fence pole traffic light traffic sign vegetation terrain sky person rider car truck bus train motorcycle bicycle Metric
DAFormer [hoyer2022daformer] 48.94 31.78 84.12 26.51 31.06 24.92 6.83 7.73 71.86 40.63 91.23 44.13 2.1 78.02 55.02 25.24 84.89 61.37 29.59 44.53
Trans4PASS (S) [zhang2022bending] 74.28 30.68 84.79 28.45 36.60 27.90 15.40 15.11 78.99 33.71 93.28 49.69 6.79 82.00 46.42 51.79 72.84 68.34 40.63 49.35
Trans4PASS+ (S) [zhang2024behind] 74.74 34.17 85.43 23.46 42.78 28.45 17.14 18.08 79.64 29.41 93.00 48.74 5.40 82.76 50.54 44.91 91.49 61.49 40.20 50.11
360SFUDA++ [zheng2024360sfuda++] 72.29 45.24 84.91 32.11 38.18 29.83 19.28 17.98 76.02 27.23 92.01 53.14 18.74 81.35 44.42 45.38 82.84 60.93 36.92 50.46
Ours 87.67 59.56 89.68 50.88 45.10 42.62 34.02 22.73 81.21 45.19 94.59 72.24 14.42 86.62 64.88 69.16 95.27 82.98 63.92 63.30
TABLE C.1: Panoramic semantic segmentation results on the DenseOoS benchmark. The per-class results are reported as IoU, and the metric is mIoU.
To provide a thorough analysis of panoramic segmentation and demonstrate that strong closed-set segmentation performance ensures effective out-of-distribution segmentation, we present the accuracy results for each inlier class in the DenseOoS dataset, as shown in Table C.1. To benchmark our model, we conduct a comparative study with the existing state-of-the-art panoramic semantic segmentation methods, namely DAFormer [hoyer2022daformer], Trans4PASS [zhang2022bending], Trans4PASS+ [zhang2024behind], and 360SFUDA++ [zheng2024360sfuda++]. DAFormer introduces domain-adaptive transformers to improve semantic segmentation across domains. Trans4PASS and its enhanced variant, Trans4PASS+, employ transformer-based architectures specifically designed to mitigate distortions and exploit contextual dependencies in panoramas. 360SFUDA++ represents an advanced source-free domain adaptation framework based on self-training, tailored to panoramic scenarios. While these methods achieve good performance under closed-set conditions, they experience notable degradation in the presence of out-of-distribution regions, owing to their limited generalization beyond the predefined semantic space. This limitation underscores the importance of enhancing robustness to unexpected content in real-world panoramic scenes.
As shown in the experimental results, our model achieves a significant mIoU of $63.30\%$, surpassing the performance of the previous best model by a large margin of $12.84\%$. A detailed breakdown of per-class performance reveals that our model excels across numerous categories, including road, sidewalk, building, wall, fence, vegetation, and terrain etc. The excellent performance in these classes demonstrates that our model is able to adapt well to the pixel distortions and background clutter introduced by the broader FoV of panoramic images, thus achieving leading segmentation performance.
C-B Additional Ablation Study
Num Iter AuPRC FPR95 mIoU
15kk 79.19 0.49 57.49
20kk 78.22 0.57 60.14
25kk 84.04 0.51 62.71
30kk 85.56 0.45 63.30
35kk 82.08 0.56 63.14
((a))
$p_{out}$ AuPRC FPR95 mIoU
0.1 86.30 0.45 62.47
0.2 85.72 0.45 62.73
0.3 85.56 0.45 63.30
0.4 81.28 0.59 62.17
0.5 80.97 0.53 62.36
((b))
Module AuPRC FPR95 mIoU
w/o 55.66 7.44 54.11
Decoders 84.25 0.50 62.38
Pixel Dec. 85.56 0.45 63.30
Trans. Dec. 46.91 5.58 57.89
((c))
TABLE C.2: Ablation study on outlier exposure fine-tuning.
C-B1 Outlier Exposure Fine-tuning
We further empirically evaluate the impact of fine-tuning parameters during the optimization of the prompt distribution space. As shown in Table LABEL:tab:ablationIter, results on the DenseOoS dataset demonstrate that the model reaches optimal performance after approximately $30k$ iterations, after which overfitting begins. For the outlier data exposure control parameter $p_{out}$, experiments reveal that when $p_{out}{>}0.3$, both closed-set performance and out-of-distribution segmentation performance deteriorate. Based on these results, we selected $0.3$ as the optimal value, balancing outlier detection with closed-set performance.
Moreover, a comparison of fine-tuning strategies (see Table LABEL:tab:ablationStr) shows that the panoramic-oriented disentanglement (Pixel Dec.) outperforms other strategies. Specifically, fine-tuning only the MLP and linear layers (w/o) or the Transformer decoder (Trans. Dec.) fails to effectively address the complexities of panoramic images. While additional fine-tuning of all decoders performs comparably in FPR95, it does not achieve optimal results on AuPRC and mIoU, decreasing segmentation accuracy. This suggests that excessive fine-tuning can introduce semantic ambiguity, compromising segmentation accuracy. These findings verify the effectiveness of our proposed fine-tuning strategy in preserving semantic consistency and enhancing out-of-distribution detection.
Features Layer AuPRC FPR95 mIoU
$f_{4}$ 1 85.56 0.45 63.30
3 71.18 0.68 57.72
6 76.46 0.64 57.33
$f_{2}$, $f_{3}$, $f_{4}$ 3 72.38 0.94 58.96
6 63.60 1.01 58.66
9 65.91 1.06 58.79
TABLE C.3: Comparison of different features in Transformer Decoder. The $f_{2}$ and $f_{3}$ are feature maps with stride $8(c_{8})$ and $16(c_{16})$ processed by DeformAttn layers, respectively.
C-B2 Features and layers in Transformer Decoder
To further analyze the most favorable settings for PanOoS, we conduct additional ablation experiments on the decoded features and the corresponding number of layers in the transformer decoder. As shown in Table C.3, the results align with previous work: decoding only semantically rich features and reducing the number of decoder layers effectively alleviates semantic confusion, which helps to better align the logit scores and results in better outlier performance [nayal2023rba]. Specialized object queries perform worse with more decoder layers. This also supports our view that, while ensuring semantic clarity, enhancing the model’s perception of panoramic pixels can significantly improve the performance of PanOoS.
Appendix D More Qualitative Results
In Fig. C.1, we visualize the PanOoS results produced by POS and other mask-transformer-based methods. Notably, POS effectively suppresses false positives in background regions, especially at the boundaries separating inlier regions, while better preserving the smoothness of the outlier map compared to other methods, despite sharing the same mask classification training paradigm.
Figure D.0: Qualitative comparisons of PanOoS. When existing pinhole-OoS methods are transferred to PanOoS, they tend to misidentify complex background regions as outliers. Although EAM [grcic2023advantages] has a low false positive rate, it incorrectly identifies bicycles as outliers and nearly overlooks the actual outliers. In contrast, Mask2Anomaly [rai2024mask2anomaly] and RbA [nayal2023rba] suffer from significant false positives in the background regions, whereas POS delivers more precise PanOoS across the $360^{\circ}$.
We showcase more visualization results of panoramic out-of-distribution segmentation of POS on DenseOoS and QuadOoS in Figs. D.1 and D.2. These examples demonstrate that POS achieves excellent performance on the PanOoS benchmark. Compared to other mask-transformer-based methods, EAM [grcic2023advantages], Mask2Anomaly [rai2024mask2anomaly], and RbA [nayal2023rba], POS reduces false positives on the ambiguous background regions and boundaries of inliers while remaining sensitive to outliers. These improvements can be observed more prominently on QuadOoS (see Fig. D.2) under defocus and motion blur scenarios.
Figure D.1: Visualization results of PanOoS on DenseOoS. POS better distinguishes outliers from inliers and produces smoother outlier maps with fewer false positives. Zoom in for a better view.
Figure D.2: Visualization results of PanOoS on QuadOoS. Uneven exposure, defocus, and motion blur caused by dense motion pose extra serious challenges to all methods. The pinhole-OoS methods have serious omissions and false positives. However, POS achieves precise segmentation of outliers. Zoom in for a better view.
Appendix E Discussion
E-A Societal impacts
In this study, we have introduced a novel task called Panoramic Out-of-distribution Segmentation (PanOoS) and established a comprehensive benchmark incorporating various well-known baseline models. We found that these baseline models exhibit limited performance on the PanOoS task, primarily due to the pixel distortions and background clutter introduced by the panoramic broader FoV. To address this issue, we have developed POS, a solution that significantly enhances the performance of the PanOoS benchmark, outperforming existing pinhole-OoS methods: pixel-wise architectures and mask-transformer-based methods, and achieving state-of-the-art results.
PanOoS promotes more comprehensive omnidirectional scene understanding. Our work has the potential to support future anomaly detection applications for assisting people with visual impairments via robotic guide dogs and wearable robotics. However, it is also possible that it can be applied in the military field, such as for drone omnidirectional perception and autonomous strikes, endangering human safety. In addition, considering that the reliability of deep learning systems for advanced driver assistance systems is crucial, it is important to note that POS may still encounter misclassifications in challenging, intense-motion scenarios, potentially leading to erroneous predictions with adverse societal implications.
E-B Future work
Although POS aims to deal with pixel distortions and background clutter in panoramic images and shows strong performance, we note that the model is not ideal for defocus and motion blur adaptation in intense motion scenarios. These issues result in performance degradation in complex dynamic environments. To address these limitations, future studies could integrate multi-modal fusion approaches to combine various types of scene information, e.g. depth and point clouds, to enhance the model’s adaptability to defocus and motion blur. Additionally, incorporating temporal information to enhance the scene perception capabilities in dynamic environments is worth investigating to further enhance PanOoS performance, with the aim of adapting to complex dynamic conditions, as well as open-set and 3D perception tasks.